OpenStudy (anonymous):

Find the area of the region bounded by x = cost, y = e^t, 0 <=t<=pi/2 and the lines y = 1 and x = 0.

OpenStudy (anonymous):

\[x=cost~~~, y=e^t~~~, 0 \le t \le \frac{ \pi }{ 2 }\] and the lines y = 1 and x = 0.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[\large A = \int\limits_a^{b} y ~dx = \int\limits_{\alpha}^{\beta} g(t) ~f'(t) dt \]

ganeshie8 (ganeshie8):

\[\large = \int\limits_{0}^{\pi/2} e^t ~(\cos t)' dt \]

ganeshie8 (ganeshie8):

doable right

OpenStudy (anonymous):

But wait, what about the y = 1?

OpenStudy (anonymous):

Shouldn't it be y-1?

ganeshie8 (ganeshie8):

Oh i missed that part

OpenStudy (anonymous):